### v2

# EndoGeDE: Generalizable Monocular Depth Estimation with Mixture of Low-Rank Experts for Diverse Endoscopic Scenes

Abstract
Self-supervised monocular depth estimation is a significant task for low-cost and efficient 3D scene perception in endoscopy. In recent years, a series of methods are proposed to address the illumination inconsistency, while certain works also focus on the generalization of the model by efficiently finetuning the foundation models. However, the variety of illumination conditions and scene features is still the primary challenges for depth estimation in endoscopic scenes. In this work, a self-supervised framework is proposed for monocular depth estimation in diverse endoscopy. Firstly, considering the diverse features in endoscopic scenes with different tissues, a novel block-wise mixture of dynamic low-rank experts is proposed to efficiently finetune the foundation model for endoscopic depth estimation. In the proposed module, based on the input feature, different experts with a small amount of trainable parameters are adaptively selected for weighted inference, from low-rank experts which are allocated based on the generalization of each block. Moreover, a novel self-supervised training framework is proposed to jointly cope with brightness inconsistency and reflectance interference. The proposed method outperforms state-of-the-art works on SCARED dataset and SimCol dataset. Furthermore, the proposed network also achieves the best generalization based on zero-shot depth estimation on C3VD, Hamlyn and SERV-CT dataset. The outstanding performance of our model is further demonstrated with 3D reconstruction and ego-motion estimation. The proposed method could contribute to accurate endoscopy for minimally invasive measurement and surgery. The evaluation codes will be released upon acceptance, while the demo videos can be found on https://endo-gede.netlify.app/.
1 Introduction
Endoscopy, as one of primary minimally invasive surgical procedures, is widely used in the medical treatment due to its minor additional trauma and short recovery period. However, due to the limited field of view and 2D vision of endoscopes, 3D scene perception has become a significant task for efficient and accurate endoscopy. Depth estimation, as one of the most common 3D vision techniques which has been applied in industrial scenes(Liu et al., 2024), underwater scenes(Ye et al., 2023), outdoor driving scenes(Feng et al., 2023; Zhu et al., 2025) and indoor scenes(Wei et al., 2024), has also been a primary 3D scene perception method in endoscopy.
Due to the high cost of obtaining depth maps (Yang et al., 2024c) and limited inference efficiency of stereo matching (Wu et al., 2025) in endoscopy, self-supervised monocular depth estimation is the main-stream method for endoscopic scenes. Although numerous self-supervised methods, such as MonoViT(Zhao et al., 2022) and LiteMono(Zhang et al., 2023a), have been proposed for monocular depth estimation for natural scenes based on deep learning, there are two critical challenges for the application of such methods in endoscopy. Firstly, illumination conditions in endoscopic scenes often causes brightness and reflectance inconsistency. Moreover, the performance and generalization of the depth estimation model will be highly limited by the variety of scene features due to different surgical tasks and tissues, which is shown in Fig. 1.
Figure 1: Challenges (left) and performance comparison (right) for depth estimation in diverse endoscopic scenes. The size of the circle denotes the number of trainable parameters in the corresponding depth estimation network.
To deal with the illumination inconsistency, Shao et al.(Shao et al., 2022) propose to utilize constraint based on the appearance flow to alleviate brightness inconsistency. Meanwhile, intrinsic image decomposition is also a useful technique to filter the interference of the light, which is also applied in endoscopy(Li et al., 2024). However, as Fig. 1 shows, the depth at some positions with high brightness on the surface will be estimated as a higher value than the ground truth, while some light spots from reflectance on the surface will highly influence the depth value at the corresponding position. With the development of foundation models for depth estimation, a series of methods are proposed to efficiently finetune the foundation model for adaptation to the endoscopic scenes, including Surgical-DINO(Cui et al., 2024a), EndoDAC(Cui et al., 2024b) and DARES(Zeinoddin et al., 2024). However, most of the existing finetuned foundation models for endoscopic depth estimation are based on Low-Rank Adaptation (LoRA)(Hu et al., 2022), the performance of which may be constrained by the feature diversity in diverse endoscopic scenes.
To this end, a novel self-supervised framework, EndoGeDE, is proposed to boost the performance and generalization of monocular depth estimation in diverse endoscopic scenes. First of all, in the proposed framework, an end-to-end training pipeline based on the integration of a novel intrinsic image registration and brightness constraint (Shao et al., 2022) is utilized to jointly deal with brightness inconsistency and light interference. Specifically, the invariance of the reflectance image to illumination changes and the strong correlations in the gradient domain(Kim et al., 2016) are considered in the proposed intrinsic image decomposition. Furthermore, a novel parameter-efficient finetuning method, Block-wise Mixture of Low-Rank Experts (BW-MoLE), is proposed to finetune the foundation model for adaptive depth estimation in endoscopy based on diverse scene features, which is built upon LoRA and Mixture of Experts (MoE). Note that the low-rank experts are allocated based on the training quality of different Transformer blocks in the encoder of foundation model. The main contribution of this work can be summarized as the following:
  1. 1.
    A novel parameter-efficient finetuning method with block-wise mixture of dynamic low-rank experts is proposed to adaptively utilize the low-rank experts for depth estimation based on diverse features, which can boost the generalization of the model.
  2. 2.
    A novel self-supervised framework for depth estimation is proposed to jointly deal with the inconsistency of brightness and reflectance in endoscopic scenes, which can benefit the depth estimation from observations with diverse illumination conditions.
  3. 3.
    The proposed method is evaluated by self-supervised depth estimation on SCARED(Allan et al., 2021) and SimCol(Rau et al., 2024) dataset, as well as zero-shot depth estimation on C3VD(Bobrow et al., 2023), Hamlyn(Recasens et al., 2021) and SERV-CT(Edwards et al., 2022) dataset. Our method achieves state-of-the-art performance compared with existing works. Moreover, the proposed method is applied to sim-to-real test(Azagra et al., 2023) and 3D reconstruction.
2 Related Work
In this section, the related works for self-supervised monocular depth estimation in endoscopy are reviewed, as well as the application of foundation models in endoscopic depth estimation.
2.1 Self-supervised Depth Estimation for Endoscopy
Based on the method proposed by Zhou et al.(Zhou et al., 2017), Turan et al. (Turan et al., 2018) proposed the first self-supervised method for depth estimation in endoscopic scenes. In the next few years, some works for self-supervised depth estimation in endoscopy are also proposed with the employment of attention mechanism. Ozyoruk et al. (Ozyoruk et al., 2021) introduced spatial attention blocks into the depth estimation network and ego-motion estimation network, while Liu et al. (Liu and Zuo, 2023) further integrated the spatial attention and channel attention in the proposed framework. In the most recent work, Yang et al. (Yang et al., 2024c) proposed a lightweight self-supervised depth estimation method for endoscopy with the combination of convolution and self-attention. Different from direct estimation from the observation, Liu et al. (Liu et al., 2023a), (Liu et al., 2023b) proposed a novel strategy to perform depth estimation based on a depth completion network, which reconstructs dense depth maps based on RGB data and corresponding sparse depth maps. Although a large amount of methods have been proposed for self-supervised depth estimation, the application of these methods in endoscopy is still limited by illumination issues in the endoscopic images.
In 2022, Shao et al. (Shao et al., 2022) proposed AF-SfMLearner to firstly tackle the brightness inconsistency using a generalized dynamic image constraint based on the appearance flow in endoscopic scenes. Based on this, Li et al. (Li et al., 2024) recently proposed a self-supervised method based on intrinsic image decomposition to deal with the challenge of reflectance on the surface of the tissue for endoscopic depth estimation. Differently, Zhou et al. (Zhou et al., 2024) proposed a method to select candidate frames with lowest temporal variance and further assign the source frame with the highest view consistency. The most recent work, MonoPCC (Wang et al., 2025), utilized a novel photometric-invariant cycle constraint to obtain cycle-warped image from the target image to solve the problem of brightness inconsistency. In this work, a novel self-supervised training strategy is proposed to jointly deal with the illumination inconsistency and the reflectance interference in diverse endoscopic scenes.
2.2 Foundation Models for Endoscopic Depth Estimation
With the development of the foundation models for depth estimation, such as DINO(Oquab et al., 2024), series of Depth Anything(Yang et al., 2024a),(Yang et al., 2024b), the application of such models in depth estimation for various scenes has gained increasing attention. For example, Yu et al.(Yu et al., 2024) utilize DINOv2 to integrate semantic and geometric features for depth estimation in outdoor scenes. Some methods for depth estimation in endoscopy based on foundation models are also proposed in recent years. Firstly, Cui et al. (Cui et al., 2024a) proposed to finetune DINO with Low-Rank Adaptation(LoRA) for depth estimation in endoscopic scenes. Based on this, a self-supervised framework, EndoDAC (Cui et al., 2024b), based on Depth Anything model (Yang et al., 2024a) with the novel parameter-efficient finetuning strategy named as dynamic-vector LoRA (DV-LoRA) and inserted convolution block is proposed to perform endoscopic depth estimation with generalizability. Meanwhile, Zeinoddin et al. (Zeinoddin et al., 2024) also proposed a different finetuning strategy named vector-LoRA which is utilized on the Depth Anything model. In this work, the foundation model for depth estimation in diverse endoscopic scenes is proposed with a novel parameter-efficient finetuning method based on the integration of MoE and LoRA, which can strengthen the adaptation of the model to different scene features.
Figure 2: The pipeline of the proposed self-supervised learning.
3 Proposed Method
3.1 Overview
The overview of the proposed method is shown as Fig. 2. The proposed framework consists of five networks, which respectively perform optical flow prediction, appearance flow prediction, depth map prediction, intrinsic image decomposition, ego-motion estimation and camera intrinsic estimation. Firstly, networks for optical flow prediction and appearance flow prediction are mainly from AF-SfMLearner(Shao et al., 2022), while ego-motion estimation and camera intrinsics estimation are based on EndoDAC(Cui et al., 2024b). The framework of depth map prediction is a finetuned encoder-decoder network based on Depth Anything(Yang et al., 2024a), which generates the corresponding depth map from single input image. Moreover, given an input observation, the reflectance image, the shading image and the soft assignment map will be generated by the proposed intrinsic image decomposition with an encoder-decoder framework.
The self-supervised learning of the framework is divided into three stages in each end-to-end training epoch to reserve the computational resources and pretend the information interference, demonstrated in Section 3.4. The intrinsic image decomposition network is firstly trained based on the consistency between the recomposed images and the raw images at the second stage, which is introduced in Section 3.3. At the final stage, the depth map prediction network is efficiently finetuned by the proposed block-wise mixture of low-rank experts in Section 3.2. Meanwhile, other networks, except the optical flow prediction network trained at the first stage, are fully trained at the final stage based on the alignments of rigid transformations using predicted ego-motion and depth map. The photometric loss is utilized to supervise the consistency of images, defined as Eq. 1.
$I^|)\mathcal{L}_{p}(\hat{I},I)=V(\beta\frac{1-SSIM(I,\hat{I})}{2}+(1-\beta$)|I-$\hat${I}|) (1)
where $V$ is the visibility mask generated from backward optical flow following (Shao et al., 2022), $\beta=0.85$.
3.2 Parameter-efficient Finetuning for Depth Map Prediction
The depth map prediction network in the proposed framework is efficiently finetuned by block-wise mixture of dynamic low-rank experts (BW-MoLE), which is shown in Fig. 3. Specifically, in each Transformer block of the network, the linear layer which generates query, value and key, as well as MLPs in the feed forward network (FFN), are finetuned by the proposed method.
3.2.1 Mixture of Low-Rank Experts
Based on DV-LoRA(Cui et al., 2024b), multiple low-rank experts are implemented to update the output based on the frozen pretrained weights. Given the input feature, the route network $\mathcal{R}$ will predict the corresponding weights of the experts network by Eq. 2, in which the noise generated by the network $\mathcal{N}$ is also introduced for robustness. The experts with top $k$ weights $\alpha_{1},\alpha_{2},...\alpha_{k}\in\Gamma_{b}=\{\alpha_{1},\alpha_{2},...,\alpha_{E_{b}}|\alpha_{i}>\alpha_{i+1}\}$ are selected to perform weighted finetuning. The finetuning process can be formulated as Eq. 3.
$\Gamma=\mathcal{R}(X)+\delta\cdot\mathcal{N}(X)$ (2)
$Y^{\prime}=Y+\sum_{i=1}^{k}\alpha_{i}V_{i}B_{i}U_{i}A_{i}X$ (3)
where $V_{i}$ and $U_{i}$ are trainable diagonal matrices of trainable weights $B$ and $A$, while $A,B$ are also trainable weights. Given the input feature map $X$, $Y$ is the output based on the frozen pretrained weights.
Figure 3: The proposed parameter-efficient finetuning of the depth map prediction network.
3.2.2 Block-wise Allocation of Experts
The weights of the network can be optimized by the smaller portion of specific features in the corresponding domain, which leads to limited generalization or transferring of the foundation model. Therefore, in this work, the number of experts is allocated based on the generalization of each block in the foundation model. To measure the generalizability, we analyze the distribution $\mathcal{D}(\lambda)$ of the eigenvalues $\lambda\in\Sigma^{2}$ based on the singular value decomposition $W=U\Sigma V$ of the extracted weights $W$ from linear layers, as Eq. 4 shows, where $\hat{\lambda}=argmax_{\lambda}(\log_{10}(\mathcal{D}(\lambda)))$. Based on the previous work(Zhou et al., 2023), the power law is utilized to fitting the probability distribution using the Hill estimator (Eq. 5) and the Fix-finger (Yang et al., 2023) method (Eq. 6).
$\rho(\lambda)=\log_{10}(\mathcal{D}(\lambda))\varpropto\lambda^{-\tau},0.95\hat{\lambda}<\lambda<1.5\hat{\lambda}$ (4)
$\tau_{j}=1+\frac{N-j}{\sum_{i=1}^{j}\log_{10}\frac{\lambda_{N-i+1}}{\lambda_{j}}},j=1,2,...,N-1$ (5)
$\tau=\tau_{t},t=\underset{t}{argmin}(\max(|1-(\frac{\Lambda_{t}}{\lambda_{t}})^{-\tau_{t}+1}-\frac{C_{t}}{N-t}|))$ (6)
where $\Lambda_{t}=[\lambda_{t},\lambda_{t+1},...,\lambda_{N}]$, $N$ is the number of eigenvalues, $C_{t}=[0,1,2,..,N-t]$.
The power law distribution with larger $\tau$ means that the distribution of the eigenvalues from pretrained weights is with the longer tail, which denotes that more weights from the smaller portion of features. Therefore, if the distribution of the eigenvalues from the pretrained weights is with larger $\tau$, more experts are needed to finetune the model for generalization. To this end, Eq. 7 is utilized to allocate the number of experts for each block $E_{b}$ based on the distributions.
$E_{b}=\lfloor\frac{\tau_{b}}{\sum_{i=1}^{B}\tau_{i}}\times E_{0}\rfloor$ (7)
where $E_{0}$ is the total number of experts.
Figure 4: The framework for intrinsic image decomposition
3.3 Self-supervised Loss for Intrinsic Image Decomposition
The framework for intrinsic image decomposition shown in Fig. 4 is designed based on the optical flow prediction and appearance flow prediction networks (Shao et al., 2022),(Li et al., 2024). Given the input image $I$, the albedo image $A$, the shading image $S$ and the assignment mask $M$ are predicted by the intrinsic image decomposition network. The albedo image represents the inherent color and texture of a surface independent of illumination conditions, while the shading image represents the light conditions. The assignment mask represents the probability of each pixel derivation from albedo, which can denote the influence of the illumination. The relationship of the raw image, the albedo image and the shading image can be defined as I=ASI=A*S.
Firstly, the raw image will be randomly augmented with $n\in(0,2)$ times higher brightness as the image $I_{n\times}$. As the albedo image won’t be influenced by illumination changes, the raw image could be recomposed by the albedo image $A_{1\times}$, $A_{n\times}$ from the raw image $I_{1\times}$ and the randomly augmented image $I_{n\times}$ and the shading image $S_{1\times}$ from the raw image, respectively. Therefore, the supervision of the recomposition is formulated as Eq. 8.
$\mathcal{L}_{rec}=\mathcal{L}_{p}(A_{1\times}*S_{1\times},I_{1\times})+\mathcal{L}_{p}(A_{n\times}*S_{1\times},I_{1\times})$ (8)
Referring to (Kim et al., 2016), there are strong correlations between the raw image and the intrinsic images at the gradient domain. Based on Retinex algorithm(Land and McCann, 1971), the decomposed image can be estimated from the raw image and the assignment mask. To this end, the photometric loss between the gradient of the estimated decomposed image and the gradient of the predicted image will be utilized to supervise the optimization of the intrinsic image decomposition, formulated as Eq. 9.
$\mathcal{L}_{ret}=\mathcal{L}_{p}(\nabla A_{1\times},\nabla I_{1\times}*M_{1\times})+\mathcal{L}_{p}(\nabla S_{1\times},\nabla I_{1\times}*(1-M_{1\times}))$ (9)
3.4 Multi-stage Self-supervised Training
Each epoch of the end-to-end training is divided into three stages, including optical flow registration, self-supervised intrinsic image decomposition and multiple transformation alignments based on the depth map. In each step, the target image $I_{t}=I_{i}$ and the corresponding source images $I_{s}\in\{I_{i+n},I_{i-n}\}$ are given to the framework for optimization.
3.4.1 First Stage: Self-supervised Optical Flow Estimation
Following (Shao et al., 2022), the optical flow prediction network is optimized based on the loss function Eq. 10, where $O_{t}$ is the forward optical flow of the source image $I_{t}$. $I^{opt}_{s\rightarrow t}$ is generated from the source image $I_{t}$ based on the forward optical flow and Spatial Transformer.
$\mathcal{L}_{1}=\mathcal{L}_{p}(I^{\prime}_{t},I^{opt}_{s\rightarrow t})+0.001|\nabla O_{t}|*e^{-\nabla|I_{t}|}$ (10)
3.4.2 Second Stage: Self-supervised Intrinsic Image Decomposition
At the second stage, the intrinsic image decomposition network is solely optimized by the loss function $\mathcal{L}_{2}$ shown as Eq. 11, consisting of the losses demonstrated in Section 3.3 based on the decomposition of the target image and the source images.
$\mathcal{L}_{2}=\mathcal{L}_{rec}+0.1\mathcal{L}_{ret}$ (11)
3.4.3 Third Stage: Multiple Transformation Alignments
[…]
Following Eq. 12, the source images $I_{t}$ and their intrinsic images $A_{t},S_{t}$ are all transformed into the pseudo images $I_{s\rightarrow t}$, $A_{s\rightarrow t}$ and $S_{s\rightarrow t}$. Based on the pseudo intrinsic images, the pseudo target images can be reconstructed and aligned with the target image $I_{t}$ by Eq. 13. Meanwhile, the photometric loss between the albedo image of the target image and the pseudo albedo image shown as Eq. 14 is utilized to supervise the rigid transformation, without influence from the illumination. Moreover, the smoothness of the depth map $D_{t}$ and the appearance flow $A_{p}$ is supervised by the loss function Eq. 15.
$\mathcal{L}_{rec}=0.02\mathcal{L}_{p}(A_{s\rightarrow t}*S_{s\rightarrow t},I_{t})+0.01\mathcal{L}_{p}(I_{s\rightarrow t},I_{t})$ (13)
$\mathcal{L}_{if}=\mathcal{L}_{p}(A_{s\rightarrow t},A_{t})$ (14)
$\mathcal{L}_{sm}=0.01|\nabla A_{p}|\cdot e^{\nabla|I_{t}-I^{opt}_{s\rightarrow t}|}+0.001|\nabla D_{t}|\cdot e^{\nabla|I_{t}|}$ (15)
Based on the above losses, the final loss function for the optimization of the networks except optical flow prediction network is defined as Eq. 16
$\mathcal{L}_{3}=\mathcal{L}_{sm}+\mathcal{L}_{rec}+0.02\mathcal{L}_{if}$ (16)
4 Experiments and Results
4.1 Implementation Details and Metrics
[…]
$Rel_{Abs}=\frac{1}{|\mathbf{D}|}\sum_{\hat{d}\in\mathbf{D}}|d-\hat{d}|/d$ (17)
$Rel_{Sq}=\frac{1}{|\mathbf{D}|}\sum_{\hat{d}\in\mathbf{D}}\frac{|d-\hat{d}|^{2}}{d}$ (18)
$RMSE=\sqrt{\frac{1}{|\mathbf{D}|}\sum_{\hat{d}\in\mathbf{D}}|d-\hat{d}|^{2}}$ (19)
$RMSE_{Log}=\sqrt{\frac{1}{|\mathbf{D}|}\sum_{\hat{d}\in\mathbf{D}}|\log d-\log\hat{d}|^{2}}$ (20)
$\delta=\frac{1}{|\mathbf{D}|}\left|\left\{\hat{d}\in\mathbf{D}|max(\frac{d}{\hat{d}},\frac{\hat{d}}{d}<1.25)\right\}\right|$ (21)
Figure 5: Allocation of experts. ’Param. of Power Law’ denotes the $\tau$ in Eq. 4
4.2 Datasets
SCARED: SCARED(Allan et al., 2021) dataset consists of 35 videos collected from porcine cadavers using a da Vinci Xi surgical robot. Following (Shao et al., 2022), this dataset is divided into 15351, 1705, and 551 frames for training, validation, and test, respectively.
Hamlyn: Hamlyn(Recasens et al., 2021) dataset contains numerous endoscopic videos of diverse robotic surgeries. In the experiments, all 92,672 endoscopic frames from the 21 rectified videos are used for zero-shot depth estimation.
SERV-CT: SERV-CT(Edwards et al., 2022) dataset consists of 8 keyframes collected from two ex-vivo porcine cadavers. In the experiments, all 32 endoscopic frames are used for zero-shot depth estimation.
SimCol: SimCol(Rau et al., 2024) dataset contains over 36,000 colonoscopic images based on the simulation in Unity3D. Following official website, the dataset is split into 28,776 and 9,009 frames for training and test, respectively.
C3VD: C3VD(Bobrow et al., 2023) dataset consists of 22 short video sequences collected from high-fidelity colon models by a clinical colonoscope. Following previous works (Paruchuri et al., 2024), 2889 frames from 8 videos are selected for zero-shot depth estimation.
EndoMapper: EndoMapper(Azagra et al., 2023) dataset consists of realistic colonoscopic sequences collected from regular medical practice. For sim-to-real evaluation, 2 frames are selected for the display of zero-shot depth estimation.
Table 1: Comparison Results of Self-supervised Depth Estimation on SCARED. The BEST results are in BOLD. $\dagger$ denotes that the foundation model is finetuned based on (Cui et al., 2024b).
Methods Backbones Year $Rel_{Abs}\downarrow$ $Rel_{Sq}\downarrow$ $RMSE\downarrow$ $RMSE_{Log}\downarrow$ $\delta\uparrow$ TP/M
AF-SfMLearner (Shao et al., 2022) CNN 2022 0.059 0.435 4.925 0.082 0.974 14.8
MonoViT (Zhao et al., 2022) ViT 2022 0.062 0.470 5.042 0.082 0.976 27.9
LiteMono (Zhang et al., 2023a) CA 2023 0.057 0.453 4.967 0.079 0.975 3.1
Surgical-DINO (Cui et al., 2024a) DN 2024 0.059 0.427 4.904 0.081 0.974 0.1
Yang et al.(Yang et al., 2024c) CA 2024 0.062 0.558 5.585 0.090 0.962 2.0
[…]
IID-SfMLearner(Li et al., 2024) CNN 2024 0.058 0.435 4.820 0.080 0.969 14.8
DVSMono (Zhou et al., 2024) MV 2024 0.055 0.410 4.797 0.078 0.975 27.0
DARES (Zeinoddin et al., 2024) DA 2024 0.052 0.356 4.483 0.073 0.980 2.9
EndoDAC(Cui et al., 2024b) DA 2024 0.052 0.362 4.464 0.072 0.979 1.6
SfM-Diffusion (Li et al., 2025) SD 2025 0.049 0.366 4.305 0.078 0.975 12.3
MonoPCC(Wang et al., 2025) MV 2025 0.051 0.349 4.488 0.072 0.983 27.0
[…]
CA: CNN+Attention, DN: DINO, MV: MonoViT, DA: Depth Anything, SD: Stable Diffusion.
4.3 Depth Estimation Results
4.3.1 Results on Realistic Data
The proposed method is evaluated on three different realistic datasets, SCARED, Hamlyn and SERV-CT. Meanwhile, our method is compared with state-of-the-art works including methods for endoscopic scenes (MonoPCC(Wang et al., 2025), EndoDAC(Cui et al., 2024b), DVSMono(Zhou et al., 2024), DARES(Zeinoddin et al., 2024), SfM-Diffusion(Li et al., 2025), IID-SfM(Li et al., 2024), Surgical-DINO(Cui et al., 2024a), AF-SfM(Shao et al., 2022) and method from (Yang et al., 2024c)) and natural scenes (MonoViT(Zhao et al., 2022), LiteMono(Zhang et al., 2023a) and Depth Anything(Yang et al., 2024a, b)). All frameworks are trained and tested on SCARED dataset at first, results of which are shown in Table. 4.2 and Fig. 6. To evaluate the generalization of the model, all models trained on SCARED dataset are utilized to perform zero-shot depth estimation on Hamlyn and SERV-CT dataset, and the experimental results are shown in Table. 2 and Fig. 7.
According the quantitative results on multiple datasets, the proposed method outperforms existing methods with lower error ($Rel$ and $RMSE$) and higher accuracy ($\delta$) in diverse endoscopic scenes. The best generalization of our method is also demonstrated based on the zero-shot evaluation. Based on qualitative results, the proposed method can provide more accurate depth estimation for the edge of tissues or instruments and smoother depth estimation for the tissue surface without interference of the light.
Table 2: Comparison Results of Zero-shot Depth Estimation on Hamlyn and SERV-CT Dataset. The BEST results are in BOLD.
Hamlyn SERV-CT
Methods $Rel_{Abs}$ $Rel_{Sq}$ RMSERMSE $RMSE_{Log}$ $\delta$ $Rel_{Abs}$ $Rel_{Sq}$ RMSERMSE $RMSE_{Log}$ $\delta$
AF-SfMLearner(Shao et al., 2022) 0.168 4.440 13.870 0.204 0.770 0.105 1.772 11.590 0.134 0.889
MonoViT (Zhao et al., 2022) 0.193 10.512 18.028 0.220 0.769 0.103 1.566 11.482 0.136 0.895
LiteMono (Zhang et al., 2023a) 0.179 6.366 15.196 0.216 0.754 0.124 2.314 13.156 0.175 0.820
Depth Anything(Yang et al., 2024a) 0.154 3.616 12.733 0.189 0.784 0.082 1.122 9.409 0.110 0.929
Depth Anything v2(Yang et al., 2024b) 0.182 4.994 15.067 0.219 0.740 0.092 1.328 10.407 0.120 0.918
IID-SfMLearner(Li et al., 2024) 0.171 4.526 14.066 0.206 0.767 0.112 1.956 12.193 0.138 0.878
DVSMono(Zhou et al., 2024) 0.143 2.956 11.905 0.181 0.796 0.095 1.412 10.811 0.127 0.907
DARES(Zeinoddin et al., 2024) 0.164 4.456 13.625 0.201 0.783 0.084 1.179 9.932 0.113 0.930
EndoDAC(Shao et al., 2022) 0.156 3.854 12.936 0.193 0.791 0.084 1.223 9.683 0.113 0.922
MonoPCC(Wang et al., 2025) 0.158 3.889 13.205 0.194 0.782 0.091 1.265 10.123 0.117 0.915
[…]
Figure 6: Qualitative results of depth estimation on SCARED dataset. The error maps denotes the relative error distribution in the area marked by the green box.
Figure 7: Qualitative results of zero-shot depth estimation on Hamlyn and SERV-CT dataset. The error maps denotes the relative error distribution in the area marked by the green box. Note that if there is no green box, the error map denotes the errors in the whole picture. The image below the input raw image is the ground truth of depth map in corresponding area.
4.3.2 Results on Simulated Data and Sim-to-Real
To further demonstrate the performance of our method, the proposed framework is trained and evaluated on simulated datasets. Firstly, our method is compared with existing methods, which are with public training codes, on SimCol dataset. Meanwhile, the models trained on SimCol dataset are compared with previous works for zero-shot depth estimation on C3VD dataset. Furthermore, the proposed model with the pretrained weights from SimCol is also utilized to perform depth estimation on samples from EndoMapper, which can display the sim-to-real performance of our work. The quantitative results are shown in Table 3, while qualitative results including sim-to-real depth estimation are shown in Fig. 8. According to the experimental results, the outstanding performance and generalization of the proposed method are further proved.
Table 3: Comparison Results of Depth Estimation on SimCol and C3VD Dataset. The BEST results are in BOLD.
SimCol (Trained) C3VD (Zero-shot)
Methods $Rel_{Abs}$ $Rel_{Sq}$ RMSERMSE $RMSE_{Log}$ $\delta$ $Rel_{Abs}$ $Rel_{Sq}$ RMSERMSE $RMSE_{Log}$ $\delta$
AF-SfMLearner(Shao et al., 2022) 0.086 0.080 0.426 0.109 0.950 0.142 1.495 7.897 0.188 0.804
IID-SfMLearner(Li et al., 2024) 0.080 0.069 0.429 0.108 0.959 0.140 1.484 7.733 0.186 0.810
EndoDAC(Cui et al., 2024b) 0.107 0.110 0.530 0.133 0.924 0.103 0.652 5.306 0.127 0.918
[…]
Figure 8: Qualitative results of depth estimation on Simulated dataset and sim-to-real test.
Table 4: Ablations on Parameter-efficient Finetuning Strategy. The best results are in red.
PEFT SCARED / SERV-CT / Hamlyn
QKV FFN $Rel_{Abs}\downarrow$ $Rel_{Sq}\downarrow$ $RMSE\downarrow$ $RMSE_{Log}\downarrow$ $\delta\uparrow$
[…]
GW-MoLE(Gao et al., 2025) 0.049 / 0.083 / 0.144 0.326 / 1.039 / 3.391 4.303 / 9.163 / 12.034 0.069 / 0.107 / 0.179 0.983 / 0.935 / 0.808
MoLE(2/3) MoLE(2/5) 0.048 / 0.083 / 0.145 0.315 / 1.081 / 3.345 4.210 / 8.948 / 12.052 0.068 / 0.105 / 0.180 0.984 / 0.936 / 0.807
MoLE(1/3) MoLE(2/5) 0.050 / 0.075 / 0.147 0.364 / 0.854 / 3.416 4.556 / 8.240 / 12.257 0.072 / 0.096 / 0.182 0.981 / 0.950 / 0.800
MoLE(2/4) MoLE(2/5) 0.051 / 0.083 / 0.153 0.348 / 1.107 / 3.672 4.428 / 9.346 / 12.721 0.072 / 0.108 / 0.189 0.980 / 0.937 / 0.792
MoLE(2/4) MoLE(2/4) 0.052 / 0.083 / 0.156 0.392 / 1.068 / 3.973 4.666 / 8.989 / 12.941 0.074 / 0.106 / 0.192 0.980 / 0.937 / 0.792
BW-MoLE $\times$ 0.051 / 0.084 / 0.147 0.364 / 1.061 / 3.363 4.511 / 9.223 / 12.212 0.073 / 0.107 / 0.182 0.979 / 0.941 / 0.799
$\times$ BW-MoLE 0.050 / 0.082 / 0.153 0.364 / 1.066 / 3.694 4.528 / 9.299 / 12.843 0.071 / 0.108 / 0.189 0.981 / 0.931 / 0.790
$\times$ MoLE(2/5) 0.055 / 0.079 / 0.156 0.476 / 1.065 / 4.079 5.119 / 9.094 / 13.062 0.081 / 0.105 / 0.192 0.977 / 0.926 / 0.793
HydraLoRA(Tian et al., 2024) 0.052 / 0.098 / 0.155 0.361 / 1.675 / 4.066 4.525 / 10.931 / 12.986 0.073 / 0.125 / 0.191 0.982 / 0.904 / 0.796
AdaLoRA(Zhang et al., 2023b) 0.054 / 0.076 / 0.144 0.438 / 0.985 / 3.057 4.949 / 8.992 / 12.075 0.078 / 0.103 / 0.179 0.978 / 0.933 / 0.801
DV-LoRA(Cui et al., 2024b) 0.050 / 0.086 / 0.151 0.334 / 1.172 / 3.639 4.342 / 9.403 / 12.469 0.071 / 0.108 / 0.186 0.983 / 0.929 / 0.798
LoRA(Hu et al., 2022) 0.049 / 0.081 / 0.150 0.336 / 1.065 / 3.681 4.381 / 9.035 / 12.508 0.070 / 0.103 / 0.185 0.983 / 0.945 / 0.800
4.4 Ablation Studies
4.4.1 Ablations on Parameter-efficient Finetuning
Ablation studies are performed to demonstrate the effect of the proposed parameter-efficient finetuning based on block-wise mixture of low-rank experts (BW-MoLE), as Table 4.3.2 shows. Firstly, the allocation of the experts are replaced with group-wise allocation (GW-MoLE(Gao et al., 2025)) and average allocation. Note that the deeper group is also allocated with more experts in GW-MoLE according to Fig. 7. The numbers of experts are set to 2, 4, 6 and 8 for each group consisting of four Transformer blocks for group-wise allocation, while the top-ks in all blocks are equal to 2. Moreover, the number of experts is set to x with the top-k setting, represented as MoLE(k/x), for average allocation. Furthermore, the finetuning of each linear layer is demonstrated based on ablation studies. At last, the finetuning based on the proposed method is also compared with the PEFT with multiple kinds of LoRA.
4.4.2 Ablations on Self-supervised Learning
Ablation studies are performed to demonstrate the effect of the proposed self-supervised training strategy, as Table 5 shows. For our training strategy, the experiments are implemented to prove the effect of the loss from augmented image, Retinex algorithm and illumination-free alignment. Moreover, the proposed model is also trained based on the existing training pipeline AF-SfMLearner (Shao et al., 2022) and IID-SfMLearner (Liu et al., 2020) for comparison.
Table 5: Ablation studies on Self-supervised Learning (SSL) pipeline. The best results are in red.
SSL Pipeline SCARED / SERV-CT / Hamlyn
$Rel_{Abs}\downarrow$ $Rel_{Sq}\downarrow$ $RMSE\downarrow$ $RMSE_{Log}\downarrow$ $\delta\uparrow$
[…]
w/o $\mathcal{L}_{if}$ 0.048 / 0.070 / 0.141 0.332 / 0.766 / 3.109 4.350 / 7.723 / 11.774 0.069 / 0.089 / 0.175 0.983 / 0.966 / 0.810
w/o $\mathcal{L}_{ret}$ 0.050 / 0.082 / 0.149 0.346 / 1.004 / 3.508 4.455 / 8.924 / 12.470 0.071 / 0.104 / 0.184 0.983 / 0.942 / 0.798
$\mathcal{L}_{rec}$ w/o $A_{n\times}$ 0.050 / 0.076 / 0.142 0.335 / 0.904 / 3.138 4.344 / 8.562 / 11.842 0.070 / 0.097/ 0.180 0.980 / 0.959 / 0.805
AF-SfMLearner(Shao et al., 2022) 0.049 / 0.083 / 0.142 0.339 / 1.071 / 3.138 4.390 / 9.086 / 11.842 0.070 / 0.107 / 0.180 0.983 / 0.932 / 0.805
IID-SfMLearner(Li et al., 2024) 0.058 / 0.092 / 0.148 0.448 / 1.342 / 3.302 4.997 / 9.924 / 12.372 0.081 / 0.118 / 0.184 0.976 / 0.913 / 0.794
4.5 3D Reconstruction Results
3D reconstruction is a significant application of depth estimation for endoscopy. Therefore, we also perform 3D scene reconstruction based on the estimated depth maps using Open3D  [^3] https://www.open3d.org/. As Fig. 9 shows, the reconstructed scenes based on the depth maps provided by our method is with more clear and accuracy texture and appearance, compared with results from (Cui et al., 2024b) and (Wang et al., 2025).
Figure 9: 3D scene reconstruction results on data from SCARED dataset.
Figure 10: Qualitative results of ego-motion estimation on SCARED dataset. Black lines denote the ground truths.
4.6 Ego-motion Estimation Results
Ego-motion estimation, as one of important part in the self-supervised framework, is also evaluated with experiments. Four sequences from the test split of SCARED dataset are selected to perform ego-motion estimation. The accuracy of ego-motion estimation is evaluated and compared based on Absolute Trajectory Error (ATE), following previous works. Based on experimental results in Table 6 and Fig. 10, the proposed method can perform the most accurate ego-motion estimation, compared with multiple existing methods including MonoPCC(Wang et al., 2025), EndoDAC(Cui et al., 2024b), IID-SfM(Li et al., 2024) and AF-SfM(Shao et al., 2022).
Table 6: Ego-motion estimation results on SCARED dataset
[…]
Methods Seq. 1 Seq. 2 Seq. 3 Seq. 4
AF-SfMLearner(Shao et al., 2022) 0.0941 0.0841 0.0742 0.0682
Depth Anything(Yang et al., 2024a) 0.0932 0.0815 0.0794 0.0696
IID-SfMLearner(Li et al., 2024) 0.0951 0.0851 0.0764 0.0682
EndoDAC(Cui et al., 2024b) 0.0936 0.0832 0.0776 0.0687
MonoPCC(Wang et al., 2025) 0.1040 0.0842 0.0781 0.0939
[…]
5 Discussion
5.1 Outstanding Accuracy and Generalization
The proposed method is compared with state-of-the-art works across different kinds of datasets, including data from different tissue during different procedures collected by different methods. The experimental results demonstrate the outstanding performance and generalization of the proposed method. Moreover, results of 3D reconstruction and ego-motion estimation further display the outstanding performance of the proposed method, compared with existing methods.
Figure 11: Visualization of the experts’ weights based on t-SNE.
5.2 Generalization on Diverse Scenes
Based on foundation models, different from previous finetuning based on LoRA, mixture of experts is firstly introduced into the depth estimation model to boost the generalization by adapting weighted inference based on diverse input features. The effect of the block-wise allocated low-rank experts is demonstrated based on a series of ablation studies, compared with various LoRA and original MoE as well. Furthermore, based on 551 test subjects from 7 kinds of scenes in SCARED dataset and 32 subjects from 2 sequences in SERV-CT dataset, the predicted weights of experts are visualized using t-SNE in Fig. 11. The visualization result demonstrates that the experts are adaptively selected based on the input features from different scenes.
Figure 12: Comparison of results from models trained via different pipelines. The layout of this figure follows Fig.7.
5.3 Generalization on Diverse Illumination
Different from previous self-supervised training, both of the interference of reflectance and the brightness inconsistency are jointly dealt with by the proposed training framework. The superiority of our pipeline is proved based on qualitative results from ablation studies and comparison with existing pipelines. Furthermore, the proposed self-supervised training strategy is compared with previous pipelines on the sample from SCARED (first two rows) and Hamlyn (the last row) dataset in Fig. 12, which demonstrates the filtering of influence from light inconsistency by our method.
5.4 Limitations & Future Works
Although the proposed method achieves state-of-the-art performance compared with existing methods on monocular depth estimation in diverse endoscopic scenes, The inference speed of our method is lower than previous works while the speed still achieves about 30fps which could meet the real-time requirement, shown in Table.7. Although the number of training parameters remains low in the proposed framework, the step-by-step training strategy and the usage of foundation model still need lots of computational resources. In the future, more efficient training pipeline and foundation model could be developed for endoscopic depth estimation. Meanwhile, multimodal visual features such as semantic feature could boost the performance of depth estimation in endoscopy.
Table 7: Comparison of Inference Speed
Methods Speed(ms/frame)
SCARED Hamlyn
Depth Anything(Yang et al., 2024a, b) 5.0 3.0
IID-SfMLearner(Li et al., 2024) 2.0 1.0
DVSMono(Zhou et al., 2024) 12.7 10.5
EndoDAC(Cui et al., 2024b) 5.7 4.0
MonoPCC(Wang et al., 2025) 12.6 10.5
[…]
6 Conclusion
In this work, a novel self-supervised framework based on a novel parameter-efficient finetuning with mixture of experts is proposed for generalizable depth estimation in diverse endoscopic scenes. The experiments on multiple datasets demonstrate the state-of-the-art performance and generalization of the proposed pipeline in diverse endoscopic scenes. This work can contribute to the engineering application of 3D perception in endoscopy for minimally invasive surgery and diagnosis. Although the proposed method achieves satisfactory performance, the lower inference speed (about 20-30 fps) and the higher computational cost (about 23GB GPU memory for training) are still limitations which need future development.